what is the simplest formula of that compound in which x-element (atomic mass = 10g/mol) is 50 persent of total mass and y-element (atomic mass 20g/mol) is 50 persent of total mass?
1
Expert's answer
2017-06-07T10:23:11-0400
Element’s name with their symbol: X and Y
Atomic mass: 10 (for X) and 20 (for Y)
Weight % in compound: 50% (for X) and 50% (for Y)
Relative number of atom: 50/10 = 5 (for X) and 50/20 = 2.5 (for Y)
Simplest atomic ratio: 5/2.5 = 2 (for X) and 2.5/2.5 = 1 (for Y)
Empirical formula: X2Y